OpenStudy (anonymous):

how do i write the cubic equation that has roots of 8 and 2i in standard form

OpenStudy (phi):

cubic means it will have x^3 in it. It will have 3 roots. it will have 2 complex roots. The complex roots will be complex conjugates. complex conjugate means if you have: a+bi (a and b are numbers, i=sqrt(-1)), then the complex conjugate is a-bi so your roots for a cubic will be: 8, 2i, -2i that means your cubic is formed by (x-8)(x-2i)(x- -2i)= (x-8)(x-2i)(x+2i) now the hard part, multiply this out.

OpenStudy (precal):

(x-2i)(x+2i) is a difference of squares \[x^2-\left( 2i \right)^2\] \[x^2-4i^2\] \[x^2-(4)(-1)\] \[x^2+4\] \[(x-8)(x^2+4) \] finish multiplying it out

OpenStudy (phi):

I left out the obvious : (x-8)(x-2i)(x+2i)=0 we want an equation when we are done

OpenStudy (precal):

Yes I know, I just thought I would help by multiplying out (x+2i)(x-2i). If they take my last step and multiply it out they will have their cubic equation. It is always good to do some work.....